# A Generalization of von Neumann’s Reduction from the Assignment Problem to Zero-Sum Games

### v2

@@ 1 Introduction
The Minimax Theorem for zero-sum games (von Neumann, 1928), which is regarded as one of von Neumann’s spectacular contributions, not only started off [^1] Von Neumann is famously quoted as saying “As far as I can see, there could be no theory of games […] without that theorem […]. I thought there was nothing worth publishing until the Minimax Theorem was proved” (Casti, 1996, p. 19). game theory (von Neumann and Morgenstern, 1947) but also had far-reaching consequences in other areas. In particular, it is equivalent to the Linear Programming (LP) Duality Theorem and it led to Yao’s Lemma (Yao, 1977), a key tool for proving a lower bound on the worst-case performance of a randomized algorithm. The equivalence mentioned above is particularly significant since it connects cornerstone problems of the two fields of game theory and optimization. Our paper provides a fresh insight into this equivalence.
The two directions of this equivalence are qualitatively different: It is straightforward to reduce a zero-sum game to a primal-dual pair of LPs whose optimal solutions yield maximin strategies of the game (Dantzig, 1951). However, the other direction is much harder, i.e., the task of solving a linear program — or deciding that the primal or the dual program is infeasible — by finding maximin strategies of a corresponding zero-sum game. This direction follows from a series of papers published over a course of more than seven decades, starting with the seminal paper by Dantzig (1951). [^2] Interestingly, Dantzig (1982) describes a memorable meeting of his with von Neumann in which the latter outlined such a correspondence. Dantzig (1982) also attributes the LP Duality Theorem to von Neumann, again as a consequence of the same meeting, even though the first written proof is due to Tucker and his students Kuhn and Gale. However, the current reduction has an unsatisfactory aspect: it does not assign distinct roles to the two players of the reduced game, as is natural from the definition of a zero-sum game, see details below.
@@
For the difficult direction, Dantzig (1951) reduced the given primal-dual pair of LPs to a symmetric zero-sum game; such a game has symmetric maximin strategies and the value of the game is zero. In this game, each player simultaneously mimics both primal and dual LPs, i.e., a strategy of the row player (or the column player) can be separated into two parts, corresponding to solutions of the primal and dual LPs. Even though this result stood for over half a century, it is in fact incomplete as observed by Adler (2013): the constructed zero-sum game can have solutions of a particular form which cannot be interpreted as optimal primal-dual pairs or certificates of infeasibility of the primal or the dual LP. Later — by adding an additional row to the zero-sum game — this reduction was extended to work in full generality, first by Adler (2013) and later, adding a certificate of infeasibility, by von Stengel (2023) and Brooks and Reny (2023); these are all strongly polynomial (in fact log-space), many-one reductions, see Section 3 for details. However, these extensions are valid only for LPs with rational numbers, hence giving the sense that the reduction is still incomplete, since the easy direction is valid for zero-sum games with real numbers (see Section 7.2 for a detailed discussion).
Next, we describe the partial resolution by von Neumann (1953). Let $(G,\mu)$ be an instance of the assignment problem on a complete bipartite graph $G$ with $n$ vertices on each side and strictly positive weights $\mu_{ij}$ on all edges $(i,j)$. The transformed game matrix $M$ is $2n\times n^{2}$, where the $2n$ rows correspond to the $2n$ vertices of $G$ and the $n^{2}$ columns correspond to the $n^{2}$ edges of $G$. Let $l$ represent edge $(i,j)$ of $G$. Then the $(k,l)^{th}$ entry of $M$ is defined as follows:
[…]
@@
The assignment problem asks for a maximum weight matching in bipartite graph $G$ under edge weights $\mu_{ij}$. In the hide-and-seek game, the row player picks one vertex and the column player picks one edge, say $e=(i,j)$, of $G$. If the vertex is one of the endpoints of the edge, then the column player pays $\frac{1}{\mu_{ij}}$ to the row player, and otherwise, no payment is made. The row player attempts to maximize his payoff and the column player minimizes the amount she has to pay. The reason for the name “hide-and-seek” should now be clear: the column player hides in an edge and the row player attempts to guess one of its endpoints. As usual, mixed strategies will be probability distributions over pure strategies.
Consider the primal-dual pair of LPs for the assignment problem in $G$: the primal LP finds a maximum weight fractional perfect matching in $G$ and the dual LP finds a minimum generalized vertex cover. The most remarkable step in the construction given above is von Neumann’s choice of the payoff matrix. Indeed, up to scaling, this matrix is exactly the constraint matrix of the LP. He shows that for this payoff matrix, the optimal solutions to the primal and dual LPs yield maximin strategies for the column and row players, respectively. But why is the payment the reciprocal of the edge weight? Our reduction generalizes the idea of using the scaled constraint matrix as the game matrix, and from this larger perspective, we are able to dispel the mystery behind von Neumann’s reduction, see Section 5.1.
A drawback of von Neumann’s reduction is that his LPs are highly specialized and have almost a “combinatorial” feel to them: His primal LP is $\max c^{\intercal}x\ \ \text{s.t. }Ax\leq\mathbf{1},\ x\geq 0$, with all entries in $c$ being strictly positive and the constraint matrix, $A$, being the incidence matrix of the underlying complete bipartite graph, i.e., all components of this matrix are zeros or ones. That still leaves a nagging doubt: can a separation of the roles of the two players be achieved if the LPs are more general, e.g., for a more general constraint vector $b$ or for data with negative entries. In this paper, we generalize von Neumann’s hide-and-seek game along two directions:
@@
A wide range of applications are covered by these two extensions. The allocation of scarce resources, including classic economic scenarios such as the diet problem (via its dual LP) or farm planning, can be modelled with LPs which typically have all components of $A$, $b$, and $c$ positive (Chvátal, 1983, Chapter 11), as required by the first extension. Additional applications are covered by the second extension, e.g., the transportation problem which is equivalent to $b$-matching, a direct generalization of the assignment problem (Schrijver, 2003, Chapter 21).
In fact, the setting of LPs with positive $b$ and $c$ corresponds to production planning scenarios and gives rise to a natural game-theoretic interpretation extending von Neumann’s hide-and-seek game, as we discuss in Section 5.2. We call the resulting game the production planning game.
We leave the open problem of giving a complete generalization. We believe this will require a substantially new approach, since we seem to have extended von Neumann’s idea of rescaling as far as was possible. Our results include:
@@
  4. 4.
    Our generalization is valid for linear programs in real numbers.
Finally, since the reduced game always has a well-defined value, achieved by maximin strategies, it offers a new perspective on LP pairs, one of which is unbounded. The game-theoretic world allows us to distinguish such LPs; this can be interpreted as a degree of (in)feasibility, see Remark 4.4.
We end our paper by discussing evidence that von Neumann possessed an implicit understanding of the notion of polynomial time.
@@
Some previous work has proposed other generalizations of von Neumann’s hide-and-seek game. However, in contrast to our paper, none of these generalizations provides a reduction between linear programs and zero-sum games.
@@ 2 Preliminaries
We start with some conventions. Whenever we write an inequality between vectors, we assume that this inequality holds componentwise. Moreover, for a vector $v$ and a matrix $M$, we denote by $v^{\intercal}$ and $M^{\intercal}$ its transposed vector and matrix, respectively. Finally, we denote by $\mathbf{1}$ and $\mathbf{0}$ the vectors or matrices consisting of $1$- and 0-entries only; their dimension will be clear from the context.
We now introduce linear programs and zero-sum games, the main objects of our study. For text book introductions to these topics, we recommend the books by Schrijver (1998) and Nisan et al. (2007). We consider linear programs of the following form. Assume that we are given $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$. We denote their components by $a_{ij}$, $b_{i}$, and $c_{j}$, respectively. These objects specify the linear program $(P)$ together with its dual $(D)$.
@@
$x$ $\geq 0$ $y$ $\geq 0$
We write $(P,D)$ to refer to the pair of linear programs $(P)$ and $(D)$. A vector $x\in\mathbb{R}^{n}$ is called (primal) feasible if it satisfies $Ax\leq b$ and $x\geq 0$, i.e., if it satisfies the constraints of $(P)$. Similarly, a vector $y\in\mathbb{R}^{m}$ is called (dual) feasible if it satisfies $A^{\intercal}y\geq c$ and $y\geq 0$, i.e., if it satisfies the constraints of $(D)$. The weak duality theorem states that if $x$ and $y$ are primal feasible and dual feasible, respectively, then $c^{\intercal}x\leq b^{\intercal}y$. This means that whenever we find feasible $x$ and $y$ such that $c^{\intercal}x=b^{\intercal}y$, then $x$ is a feasible vector maximizing $c^{\intercal}x$ and $y$ is a feasible vector minimizing $b^{\intercal}y$. In this case, we say that $(x,y)$ an optimal primal-dual pair. The strong duality theorem states that whenever $(P)$ and $(D)$ admit feasible points, then they admit an optimal primal-dual pair, and if so, $c^{\intercal}x=b^{\intercal}y$ is called the value of the LP. The following characterization of possible outcomes of a pair of primal and dual programs $(P,D)$ is well-known.
@@
In order to certify $(P)$ to be unbounded, we need to show, in addition to the unboundedness certificate, a feasible solution to $(P)$ (i.e., a vector $x$ such that $Ax\leq b$, $x\geq\mathbf{0})$.
If $z^{*}$ is an infeasibility certificate for $(P)$, then $z^{*}$ is an unboundedness certificate for $(D)$.
If $w^{*}$ is an unboundedness certificate for $(P)$, then $w^{*}$ is an infeasibility certificate for $(D)$.
Cases (2) and (3) in Proposition 2.1 establish that the linear programs have no optimal solutions by certifying that either $(P)$ or $(D)$ have no feasible solution.
Our second important concept is zero-sum games. A zero-sum game is specified by a matrix $M=(m_{ij})\in\mathbb{R}^{m\times n}$. There are two players, a row player and a column player; for ease of notation, we will assume that the first is male and the second is female. The row player’s action is to select a row $i$ with $1\leq i\leq m$ and the column player’s action is to select a column $j$ with $1\leq j\leq n$. Selecting row $i$ and column $j$ leads to a payment of the associated matrix entry, $m_{ij}$, from the column player to the row player. Hence, whereas the row player’s goal is to maximize the value of the matrix entry chosen, the column player’s goal is to minimize it.
The players are allowed to randomize their actions. A mixed strategy is a probability distribution over actions. The row and column players select mixed strategies over rows and columns, respectively. These can be specified by vectors $p\in\mathbb{R}^{m}$ and $q\in\mathbb{R}^{n}$ that are nonnegative and satisfy $p^{\intercal}\mathbf{1}=q^{\intercal}\mathbf{1}=1$. The entries of $p$ and $q$ are the probabilities with which the respective rows and columns are played. If the row player plays $p$ and the column player plays $q$, then the row player receives the expected payment of $p^{\intercal}Mq$ from the column player.
@@
The unique payments determined by these strategies, i.e., $u^{*}=\max_{p}\min_{q}p^{\intercal}Mq$ and $v^{*}=\min_{q}\max_{p}p^{\intercal}Mq$ are called the maximin value and minimax value of the game, respectively. The famous Minimax Theorem by von Neumann (1928) states that
@@
i.e., the maximin value of the game coincides with its minimax value. Therefore, this quantity is simply called the value of the game. The goal of solving a zero-sum game is to compute a pair of maximin strategies. Note that such strategies are mutual best responses, and therefore constitute a Nash equilibrium.
It is well known that maximin strategies as well as the value of the game can be computed by the following linear programs; this is the easy direction of the correspondence between linear programs and zero-sum games (see, e.g., von Stengel, 2023, Equations (7) and (8)):
@@
s.t. $M^{\intercal}p\geq\mathbf{1}u$ s.t. $Mq\leq\mathbf{1}v$
@@
Note that $(\mbox{{Row}},\mbox{{Col}})$ are a primal-dual pair. The optimal value of the linear program $(\mbox{{Row}})$ is the maximin value corresponding to the row player. An optimal solution $(u^{*},p^{*})$ of $(\mbox{{Row}})$ consists of the maximin value together with a maximin strategy. In fact, the constraints of $(\mbox{{Row}})$ imply that $p$ is a probability distribution and that $p^{\intercal}M\geq\mathbf{1}^{\intercal}u$. Therefore, ${p^{*}}^{\intercal}Mq\geq u^{*}$ for all mixed strategies $q$ of the column player. Similarly, $(\mbox{{Col}})$ determines the minimax value and a minimax strategy of the column player.
3 Reducing Linear Programs to Zero-Sum Games
In this section, we lay the groundwork for performing the difficult direction of the correspondence, i.e., reductions from LPs to zero-sum games. We start by reviewing the ideas behind the original reduction by Dantzig (1951).
In his seminal paper already discussed in the introduction, Dantzig (1951) introduced a reduction from LPs to zero-sum games by using a skew-symmetric game matrix. More precisely, given the primal LP $(P)$, he considers the zero-sum game with payoff matrix [^3] Dantzig (1951) attributes the idea for this matrix to Tucker.
As this matrix is skew-symmetric (so it is a symmetric game), the value of the game is 0. Denoting a maximin strategy as $(p^{*},q^{*},t^{*})$, the key result of the paper by Dantzig (1951) states that
  1. 1.

### v3

@@ 1 Introduction
The Minimax Theorem for zero-sum games (von Neumann, 1928), which is regarded as one of von Neumann’s spectacular contributions, not only started off [^1] Von Neumann is famously quoted as saying “As far as I can see, there could be no theory of games […] without that theorem […]. I thought there was nothing worth publishing until the Minimax Theorem was proved” (Casti, 1996, p. 19). game theory (von Neumann and Morgenstern, 1947) but also had far-reaching consequences in other areas. In particular, it is equivalent to the Linear Programming Duality Theorem and it led to Yao’s Lemma (Yao, 1977), a key tool for proving a lower bound on the worst-case performance of a randomized algorithm. The equivalence mentioned above is particularly significant since it connects cornerstone problems of the two fields of game theory and optimization. Our paper provides a fresh insight into this equivalence.
The two directions of this equivalence are qualitatively different: It is straightforward to reduce a zero-sum game to a primal-dual pair of linear programs (LPs) whose optimal solutions yield maximin strategies of the game (Dantzig, 1951). However, the other direction is much harder, i.e., the task of solving a linear program — or deciding that the primal or the dual program is infeasible — by finding maximin strategies of a corresponding zero-sum game. This direction follows from a series of papers published over a course of more than seven decades, starting with the seminal paper by Dantzig (1951). [^2] Interestingly, Dantzig (1982) describes a memorable meeting of his with von Neumann in which the latter outlined such a correspondence. Dantzig (1982) also attributes the LP Duality Theorem to von Neumann, again as a consequence of the same meeting, even though the first written proof is due to Tucker and his students Kuhn and Gale. However, the current reduction has an unsatisfactory aspect: it does not assign distinct roles to the two players of the reduced game, as is natural from the definition of a zero-sum game, see details below.
@@
For the difficult direction, Dantzig (1951) reduced the given primal-dual pair of LPs to a symmetric zero-sum game; such a game has symmetric maximin strategies and the value of the game is zero. In this game, each player simultaneously mimics both primal and dual LPs, i.e., a strategy of the row player (or the column player) can be separated into two parts, corresponding to solutions of the primal and dual LPs. Even though this result stood for over half a century, it is in fact incomplete as observed by Adler (2013): the constructed zero-sum game can have solutions of a particular form which cannot be interpreted as optimal primal-dual pairs or certificates of infeasibility of the primal or the dual LP. Later — by adding an additional row to the zero-sum game — this reduction was extended to work in full generality, first by Adler (2013) and later, adding a certificate of infeasibility, by von Stengel (2023) and Brooks and Reny (2023).
Next, we describe the partial resolution by von Neumann (1953). Let $(G,\phi)$ be an instance of the assignment problem on a complete bipartite graph $G$ with $n$ vertices on each side and a strictly positive weight $\phi_{ij}$ on edge $(i,j)$. The transformed game matrix $M$ is $2n\times n^{2}$, where the $2n$ rows correspond to the $2n$ vertices of $G$ and the $n^{2}$ columns correspond to the $n^{2}$ edges of $G$. Let $\ell$ represent edge $(i,j)$ of $G$. Then the $(k,\ell)^{th}$ entry of $M$ is defined as follows:
$M(k,\ell)\ \ =\ \begin{cases}1/\phi_{ij}&\text{if }k=i\ \text{ or }k=j,\\ 0&\text{otherwise}.\end{cases}$
@@
The assignment problem asks for a maximum weight matching in bipartite graph $G$ under edge weights $\phi_{ij}$. In the hide-and-seek game, the row player picks one vertex and the column player picks one edge, say $e=(i,j)$, of $G$. If the vertex is one of the endpoints of the edge, then the column player pays $\frac{1}{\phi_{ij}}$ to the row player, and otherwise, no payment is made. The row player attempts to maximize his payoff and the column player minimizes the amount she has to pay. The reason for the name “hide-and-seek” should now be clear: the column player hides in an edge and the row player attempts to guess one of its endpoints. As usual, mixed strategies will be probability distributions over pure strategies.
Consider the primal-dual pair of LPs for the assignment problem in $G$: the primal LP finds a maximum weight fractional perfect matching in $G$ and the dual LP finds a minimum generalized vertex cover. The most remarkable step in the construction given above is von Neumann’s choice of the payoff matrix. Indeed, up to a change in variables caused by scaling, this matrix is exactly the constraint matrix of the LP. He shows that for this payoff matrix, the optimal solutions to the primal and dual LPs yield maximin strategies for the column and row players, respectively. But why is the payment the reciprocal of the edge weight? Our reduction generalizes the idea of using the scaled constraint matrix as the game matrix. From this larger perspective, we are able to dispel the mystery behind von Neumann’s reduction, see Section 4.1.
A drawback of von Neumann’s reduction is that his LPs are highly specialized and have almost a “combinatorial” feel to them: His primal LP is $\max c^{\intercal}x\ \ \text{s.t. }Ax\leq\mathbf{1},\ x\geq\mathbf{0}$, with all entries in $c$ being strictly positive and the constraint matrix, $A$, being the incidence matrix of the underlying complete bipartite graph, i.e., all components of this matrix are zeros or ones. That still leaves a nagging doubt: can a separation of the roles of the two players be achieved if the LPs are more general, e.g., for a more general constraint vector $b$ or for data with negative entries. In this paper, we generalize von Neumann’s hide-and-seek game along two directions:
@@
A wide range of applications are covered by these two extensions. The allocation of scarce resources, including classic economic scenarios such as the diet problem (via its dual LP) or farm planning, can be modelled with LPs which typically have all components of $A$, $b$, and $c$ positive (Chvátal, 1983, Chapter 11), as required by the first extension. Additional applications are covered by the second extension, e.g., the transportation problem which is equivalent to $b$-matching, a direct generalization of the assignment problem (Schrijver, 2003, Chapter 21). In fact, the setting of LPs with positive $b$ and $c$ corresponds to production planning scenarios and gives rise to a natural game-theoretic interpretation extending von Neumann’s hide-and-seek game, as we discuss in Section 4.1. We call the resulting game the production planning game.
We leave the open problem of giving a complete generalization. We believe this will require a substantially new approach, since we seem to have fully extended von Neumann’s idea of rescaling. Our results include:
@@
Finally, since the reduced game always has a well-defined value, achieved by maximin strategies, it offers a new perspective on LP pairs, one of which is unbounded. The game-theoretic world allows us to distinguish such LPs; this can be interpreted as a degree of (in)feasibility, see Section 3.1.
@@
Some previous work has proposed other generalizations of von Neumann’s hide-and-seek game. However, in contrast to our paper, none of these provides a reduction between linear programs and zero-sum games.
@@ 2 Preliminaries
We start with some conventions. Whenever we have an inequality between vectors, we assume that it holds componentwise. Moreover, for a vector $v$ and a matrix $M$, we denote by $v^{\intercal}$ and $M^{\intercal}$ its transposed vector and matrix, respectively. Finally, we denote by $\mathbf{1}$ and $\mathbf{0}$ the column vectors or matrices consisting of $1$- and 0-entries only; their dimension will be clear from the context.
We now introduce linear programs and zero-sum games, the main objects of our study. For text book introductions to these topics, we recommend the books by Schrijver (1998) and Nisan et al. (2007).
We consider linear programs of the following form. Assume we are given $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$. We denote their components by $a_{ij}$, $b_{i}$, and $c_{j}$, respectively. These objects specify the LP $(P)$ together with its dual $(D)$.
@@
$x$ $\geq\mathbf{0}$ $y$ $\geq\mathbf{0}$
We write $(P,D)$ to refer to the pair of LPs $(P)$ and $(D)$. A vector $x\in\mathbb{R}^{n}$ is called (primal) feasible if it satisfies $Ax\leq b$ and $x\geq\mathbf{0}$, i.e., if it satisfies the constraints of $(P)$. Similarly, a vector $y\in\mathbb{R}^{m}$ is called (dual) feasible if it satisfies $A^{\intercal}y\geq c$ and $y\geq\mathbf{0}$, i.e., if it satisfies the constraints of $(D)$. The weak duality theorem states that if $x$ and $y$ are primal feasible and dual feasible, respectively, then $c^{\intercal}x\leq b^{\intercal}y$. This implies that whenever we find feasible $x$ and $y$ such that $c^{\intercal}x=b^{\intercal}y$, then $x$ is a feasible vector maximizing $c^{\intercal}x$ and $y$ is a feasible vector minimizing $b^{\intercal}y$. In this case, we say that $(x,y)$ an optimal primal-dual pair. The strong duality theorem states that whenever $(P)$ and $(D)$ admit feasible points, then they admit an optimal primal-dual pair, and if so, $c^{\intercal}x=b^{\intercal}y$; this is called the value of the LP. The following characterization of possible outcomes of a pair of primal and dual LPs $(P,D)$ is well-known.
@@
$(i)$ In order to certify $(P)$ to be unbounded, we need to show, in addition to the unboundedness certificate, a feasible solution to $(P)$ (i.e., a vector $x$ such that $Ax\leq b$, $x\geq\mathbf{0})$.
$(ii)$ If $z^{*}$ is an infeasibility certificate for $(P)$, then $z^{*}$ is an unboundedness certificate for $(D)$. Similarly, if $w^{*}$ is an unboundedness certificate for $(P)$, then $w^{*}$ is an infeasibility certificate for $(D)$.
$(iii)$ Cases (2) and (3) in Proposition 2.1 establish that the LPs have no optimal solutions by certifying that either $(P)$ or $(D)$ are infeasible.
Our second important concept is zero-sum games. A zero-sum game is specified by a matrix $M=(m_{ij})\in\mathbb{R}^{m\times n}$. There are two players, a row player and a column player; for ease of description, we will assume that the first is male and the second is female. The row player’s action is to select a row $i$ with $1\leq i\leq m$ and the column player’s action is to select a column $j$ with $1\leq j\leq n$. Selecting row $i$ and column $j$ leads to a payment of the associated matrix entry, $m_{ij}$, from the column player to the row player. Hence, whereas the row player’s goal is to maximize the value of the matrix entry chosen, the column player’s goal is to minimize it.
Players are allowed to randomize their actions. A mixed strategy is a probability distribution over actions. The row and column players select mixed strategies over rows and columns, respectively. These can be specified by vectors $p\in\mathbb{R}^{m}$ and $q\in\mathbb{R}^{n}$ that are nonnegative and satisfy $p^{\intercal}\mathbf{1}=q^{\intercal}\mathbf{1}=1$. The entries of $p$ and $q$ are the probabilities with which the respective rows and columns are played. If the row player plays $p$ and the column player plays $q$, then the row player receives the expected payment of $p^{\intercal}Mq$ from the column player.
@@
The unique payments determined by these strategies, i.e., $u^{*}=\max_{p}\min_{q}p^{\intercal}Mq$ and $v^{*}=\min_{q}\max_{p}p^{\intercal}Mq$ are called the maximin value and minimax value of the game, respectively. The famous Minimax Theorem of von Neumann (1928) states that
@@
i.e., the maximin value of the game coincides with its minimax value. Therefore, this quantity is called the value of the game. The goal of solving a zero-sum game is to compute a pair of maximin strategies. Note that such strategies are mutual best responses, and therefore constitute a Nash equilibrium.
It is well known that maximin strategies as well as the value of the game can be computed by the following LPs; this is the easy direction of the correspondence between LPs and zero-sum games (see, e.g., von Stengel, 2023, Equations (7) and (8)):
@@
s.t. $M^{\intercal}p\geq u\mathbf{1}$ s.t. $Mq\leq v\mathbf{1}$
@@
Note that $(\mbox{{Row}},\mbox{{Col}})$ are a primal-dual LP pair. The optimal value of the linear program $(\mbox{{Row}})$ is the maximin value corresponding to the row player. An optimal solution $(u^{*},p^{*})$ of $(\mbox{{Row}})$ consists of the maximin value together with a maximin strategy. In fact, the constraints of $(\mbox{{Row}})$ imply that $p$ is a probability distribution and that $p^{\intercal}M\geq\mathbf{1}^{\intercal}u$. Therefore, ${p^{*}}^{\intercal}Mq\geq u^{*}$ for all mixed strategies $q$ of the column player. Similarly, $(\mbox{{Col}})$ determines the minimax value and a minimax strategy of the column player.
3 Extending Von Neumann’s Reduction
In this section, we provide our extensions of von Neumann’s reduction. First, we consider the case where $b>\mathbf{0}$ and $c>\mathbf{0}$, i.e., both vectors are strictly positive. Subsequently, we will show how this enables us to tackle the case that the constraint matrix is nonnegative, i.e., $A\geq\mathbf{0}$.
3.1 Positive Objective and Constraint Vectors
The key technical step of our reduction is a rescaling of the LPs $(P)$ and $(D)$; this is where we require $b>\mathbf{0}$ and $c>\mathbf{0}$. Consider the diagonal matrices $B\in\mathbb{R}^{m\times m}$ and $C\in\mathbb{R}^{n\times n}$ defined by